OpenStudy (anonymous):

Determine the factors of 4x^2 + 5x - 6. (4x - 2)(x + 3) (4x - 3)(x + 2) (4x - 6)(x + 1) (4x - 1)(x + 6)

OpenStudy (anonymous):

@Cosmichaotic @jim_thompson5910 @lasttccasey

OpenStudy (lasttccasey):

(4x - 3)(x + 2)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (lasttccasey):

You're welcome. It's rather hard to explain these over the internet but sure.

OpenStudy (anonymous):

Determine the factors of 4ry + 3rz + 8xy + 6xz. (4x + 4)(2y + 3z) (4x + 3z)(2y + r) (4y + r)(2x + 3z) (4y + 3z)(2x + r)

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

Isn't it D?

OpenStudy (lasttccasey):

(4*y + 3*z)*(r + 2*x) so yes D

OpenStudy (lasttccasey):

What class is this for if you don't mind me asking?

OpenStudy (anonymous):

Okay great! Its for Algebra.

OpenStudy (anonymous):

Factor completely: 36x^2 - 49 (3x + 7)(12x - 7) (3x - 7)(12x + 7) (6x - 7)(6x - 7) (6x + 7)(6x - 7)

OpenStudy (lasttccasey):

Well the fastest, somewhat cheating, method is to simply FOIL all the given choices. This backfires when you aren't given multiple choice tests though.

OpenStudy (anonymous):

This one would also be D right?

OpenStudy (lasttccasey):

Since this one have two perfect square roots its rather easy to find but it becomes (6*x - 7)*(6*x + 7) so yeah D

OpenStudy (anonymous):

Which polynomial is a difference of two squares? x^2 + 25 x^2 - 25 x^2 - 10 x^2 + 10 Is this one C? I don't quite understand how to do it.

OpenStudy (lasttccasey):

This page gives a good explanation http://www.purplemath.com/modules/specfact.htm

OpenStudy (lasttccasey):

x^2 - 25 is the answer to this one

OpenStudy (anonymous):

Thanks!

OpenStudy (lasttccasey):

You're welcome bud.

OpenStudy (anonymous):

What are the zeroes of f(x) = x^2 + 5x + 6? x = -2, -3 x = 2, 3 x = -2, 3 x = 2, -3